an astronaut weighs 99N on the moon, where the acceleration of gravity is 1.62 m / s2. How much does the astronaut weigh on the earth?
Answer #1

We know that

W=mg

99=m*1.62

m=61.11kg

So we have mass = 61.11 kg

And we know that value of g = 9.8 in earth

So

weight on earth =mg =61.11*9.8=598.89N